OpenStudy (anonymous):

What is the sum of the first five terms of the series 35+30+25+...?

OpenStudy (anonymous):

35+30+25+20+15=125

OpenStudy (shamil98):

You can see each consecutive term has a difference of 5, right? Your first term is 35. Using the sum formula for an arithmetic series: \[\frac{ n }{ 2 }(2a +(n-1)d)\] a = first term n = number of terms d = difference between terms You could just count it like above but it's nice to know the formula for when you are asked to find the sum of a much higher number of terms and it becomes more tedious to do by hand :)

OpenStudy (anonymous):

Thank you to both of you :3 Very much help^^